OpenStudy (anonymous):

Use the formula for the sum of the first n terms of a geometric sequence to solve. Find the sum of the first 8 terms of the geometric sequence: -8, -16, -32, -64, -128, . . . . A. -2003 B. -2040 C. -2060 D. -2038

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (anonymous):

i thought it was b

OpenStudy (michele_laino):

the first term is: a1=-8, the constant is d=2 so the requested sum, is: \[\Large {S_8} = {a_1}\frac{{1 - {q^8}}}{{1 - q}} = - 8 \cdot \frac{{1 - 256}}{{1 - 2}} = ...?\]

OpenStudy (anonymous):

\[r=\frac{ -16 }{ -8 }=2\]

OpenStudy (anonymous):

i got the answer thnx for the help doe

OpenStudy (anonymous):

Find the probability. What is the probability that a card drawn from a deck of 52 cards is not a 10?

OpenStudy (anonymous):

i need help with this one though

OpenStudy (anonymous):

A. 12/13 B. 9/10 C. 1/13 D. 1/10 is my answers

OpenStudy (anonymous):

\[\frac{ C_{1}^{48} }{ C _{1}^{52} }=\frac{ 48 }{ 52 }=?\]

OpenStudy (anonymous):

im not sure

OpenStudy (anonymous):

@Nnesha @dan815

OpenStudy (anonymous):

@Mehek14

OpenStudy (anonymous):

it would be 1/13 right

OpenStudy (anonymous):

\[\frac{ 48 }{ 52 }=\frac{ 12 }{ 13 }\]